# DiffImp: Efficient Diffusion Model for Probabilistic Time Series Imputation with Bidirectional Mamba Backbone

### v1

@@ Abstract
Probabilistic time series imputation has been widely applied in real-world scenarios due to its ability to estimate uncertainty of imputation results. Meanwhile, denoising diffusion probabilistic models (DDPMs) have achieved great success in probabilistic time series imputation tasks with its power to model complex distributions. However, current DDPM-based probabilistic time series imputation methodologies are confronted with two types of challenges: 1)  The backbone modules of the denoising parts are not capable of achieving sequence modeling with low time complexity. 2) The architecture of denoising modules can not handle the inter-variable and bidirectional dependencies in the time series imputation problem effectively. To address the first challenge, we integrate the computational efficient state space model, namely Mamba, as the backbone denosing module for DDPMs. To tackle the second challenge, we carefully devise several SSM-based blocks for bidirectional modeling and inter-variable relation understanding. Experimental results demonstrate that our approach can achieve state-of-the-art time series imputation results on multiple datasets, different missing scenarios and missing ratios.
1 Introduction
@@
The analysis of time series can model the intrinsic patterns within time-series data, thus providing robust support for decision-making in various fields, such as meteorology McGovern et al. (2011); Karevan & Suykens (2020), financial analysis Xiang et al. (2022); Owusu et al. (2023); Bai et al. (2020), healthcare Morid et al. (2023); Poyraz & Marttinen (2023) and power systems Tzelepi et al. (2023); Zhou et al. (2021). To enhance the reliability of analytical outcomes, it is critical to ensure the integrity of time series. However, due to various reasons such as device failures, human errors, and privacy protection, time series data can easily be incomplete with missing observations at different timestamps.
@@
Time series imputation methods aim to estimate the values of missing points based on the observed points in incomplete time series, thereby restoring the integrity of the time series while preserving its original statistical properties. According to the ability to provide uncertainty of estimations, time series imputation methods can be categorized into the following two perspectives: 1) Deterministic Cao et al. (2018); Cini et al. (2022); Du et al. (2023), and 2)  Probabilistic Chen et al. (2023); Kim et al. (2023); Luo et al. (2018) imputation methods. Probabilistic time series imputation is particularly important in dealing with complex and uncertain data environments, as it provides a quantification of uncertainty for the imputations. The key to probabilistic imputation lies in modeling the posterior distribution. Existing probabilistic time series imputation methods include Gaussian Process and Variational Autoencoder-based methods Fortuin et al. (2020), Normalization Flow-based methods Rasul et al. (2021), and Diffusion-based methods Tashiro et al. (2021). Among these, the Diffusion-based method has emerged as the optimal choice for probabilistic time series due to their accuracy in posterior modeling and adaptability to different scenarios and various types of time series data.
When selecting a denoising backbone in the diffusion model, the following two key factors need to be considered: 1) Model compatibility, and 2) Time complexity. Specifically, the missing observations in time series have correlations with their neighbors on both sides, so it is crucial to design a model by considering information from neighbors of both sides. Moreover, it is also essential to accurately capture the properties of time series, such as global dependencies and channel correlations. Three mainstream denoising backbones are widely used in diffusion models for time series imputation: 1) Convoluational Neural Networks (CNNs)-, 2) Transformer- and 3) State-Space Model (SSM)-based backbones. Given a time series with a length of $L$, the CNNs-based backbone can capture partial information from the neighbors within the receptive fields and has $\mathcal{O}(L)$ time complexity. The transformer-based backbone can model temporal dependencies across the entire time series but is with quadratic time complexity $\mathcal{O}(L^{2})$. The SSM backbone has a linear time complexity, $\mathcal{O}(L)$, but it falls short in capturing the information from one side of the neighbor. Moreover, all these backbones fail to capture the channel dependencies in time series. The comparison results of existing backbones and our method in terms of various dependencies and time complexity are presented in Table.1.
@@
Table 1: Comparison of our method and existing methods in modeling dependencies and time complexity. The results show that our method achieves the most comprehensive data modeling with the lowest time complexity.
Backbone Model Global Dependency Time Complexity Channel Dependency Inter-sequence Dependency
CNN Local $\mathcal{O}(L)$ Independent Unidirectional
Transformer Global $\mathcal{O}(L^{2})$ Independent Unidirectional
SSM Partial $\mathcal{O}(L)$ Independent Unidirectional
DiffImp (Ours) Global $\mathcal{O}(L)$ Dependent Bidirectional
In this paper, we propose an efficient diffusion-based framework for probabilistic time series imputation to address the drawbacks in existing backbones of time series imputation, we name it DiffImp. To ensure linear complexity, we choose the SSM-based model as the backbone of our framework, which is Mamba Dao & Gu (2024) to be more specific. To enable Mamba to capture information from both sides of the missing values, we then propose a Bidirectional Attention Mamba block (BAM) that is more applicable to time series imputation task. To incorporate bidirectional dependencies, we design a learnable weight module inside the BAM block. This module learns the weights of all points within the sequence, facilitating the modeling of dependencies at different distances. Next, we propose a Channel Mamba Block (CMB) to capture the dependencies among different channels in a time series. Specifically, we treat the variables across different channels in the time series as a sequence of variables and employ the Mamba model alongside the channel dimension, so inter-dependencies among channels can be modeled.
@@
  • We propose DiffImp, an efficient diffusion-based model for the time series imputation task. We leverage the advantages of the bidirectional structure to devise blocks that can model sequence information and channel dependencies as the denoising backbone for the diffusion modules for performance improvements.
  • We dive deeply into the application of Mamba blocks by exploring the application of Mamba as the backbone block of the DDPMs for probabilistic time series imputation. We integrate the Mamba-based blocks to the diffusion models, therefore achieving effectively time series modeling with linear time complexity.
  • We conduct experiments on multiple real-world datasets for both time series imputation and time series forecasting tasks. The experimental results demonstrate that our approach achieves state-of-the-art performance across several datasets, different missing scenarios and missing ratios.
2 Preliminaries
2.1 State Space Models
State Space Models (SSMs) are an emerging approach to model sequential data, which is implemented by finding out state representations to model the relationship between input and output sequences. A SSM receives a one-dimensional sequence $X\in\mathbb{R}^{L}$ as the input and outputs a corresponding sequence $Y\in\mathbb{R}^{M}$. Under continuous settings, the SSMs are defined according to Eq.1:
$\begin{cases}\dot{h}(t)&=\bm{A}h(t)+\bm{B}x(t)\\ y(t)&=\bm{C}h(t)+\bm{D}x(t),\end{cases}$ (1)
@@ 2.1 State Space Models
In real-world applications, the input sequences are discrete samplings of continuous sequences. According to  Gu et al. (2022), under discrete settings, by applying the zero-order hold technique to Eq.1, it can be reformulated as follows.
$\begin{cases}h_{k}=\bm{\bar{A}}h_{k-1}+\bm{\bar{B}}x_{k}\\ y_{k}=\bm{C}h_{k}\end{cases},$ (2)
where $\bm{\bar{A}}=\exp(\Delta\bm{A}),\bm{\bar{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot(\Delta\bm{B})$ and $\Delta$ is the learnable step size in discrete sampling. We can see from Eq.2 that the hidden state is updated according to the input $x(t)$ and last hidden state $h(t-1)$ while the output is generated by the hidden state $h(t)$ and the input $x(t)$ and in Gu et al. (2020), where it introduces High-order Polynomial Projection Operator (Hippo) to achieve longer sequence modeling.
However, it is worth noticing that $\bm{A},\bm{B},\bm{C},\bm{D}$ in Eq.1 and Eq.2 are time-invariant parameters, i.e., they are data-independent parameters and do not change over time. Therefore the model is not capable of assigning different weights at different positions in the input sequence while receiving new inputs. To address this issue, Gu & Dao (2023) proposed Mamba, in which the parameter matrices $\bm{A},\bm{B},\bm{C},\bm{D}$ are input-dependent, thus enhancing the performance of sequence modeling. To tackle the problem of non-parallelization, Gu & Dao (2023) also introduced selective scan mechanism for effective computing. For further performance and efficiency improvements, Dao & Gu (2024) point out that SSMs can be categorized as a variant of linear attention model. In this work, we follow the same architecture of parallel Mamba Blocks as Dao & Gu (2024) and a RMS-norm Zhang & Sennrich (2019) module is added after the parallel Mamba block. The details of the post-normalization Mamba Block (PNM Block) are illustrated in Fig.3a.
2.2 Diffusion Models
@@
$p_{\theta}(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),$ (3)
@@
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}(x_{t},t)I),$ (4)
The loss function of DDPM aims at minimizing the difference between the noise in the forward process $\epsilon$ and the parameterized noise $\epsilon_{\theta}$ in the reverse process:
$\mathcal{L}_{d}=\mathbb{E}_{x_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|,$ (5)
where $t$ stands for the diffusion time embedding and $x_{t}$ is calculated in the forward process. Please refer to Appendix 7.1 for more details about the diffusion models.
2.3 Problem Formulation
Definition 1 (Time Series).
@@ 2.3 Problem Formulation
Definition 2 (Probabilistic Time Series Imputation).
Given an incomplete time series $\tilde{X}=(X,M,T)$, where $\sum M<K\cdot L$, the problem of probabilistic time series imputation is to learn an imputation function $\mathcal{M}_{\theta}$, such that
$\bar{X}=\mathcal{M}_{\theta}(\tilde{X}),$ (6)
@@
where $\bar{X}\in\mathbb{R}^{K\times L}$ is the imputed time series, where $\bar{X}_{i,j}=\mu_{i,j}\pm\sigma_{i,j}$ denotes the probabilistic output if $M_{i,j}=1$, otherwise $\bar{X}_{i,j}={X}_{i,j}$.
3 Methodology
3.1 Diffusion Models for Time Series Imputation

### v2

# SSD-TS: Exploring the Potential of Linear State Space Models for Diffusion Models in Time Series Imputation

@@ Abstract.
Probabilistic time series imputation has been widely applied in real-world scenarios due to its ability for uncertainty estimation and denoising diffusion probabilistic models (DDPMs) have achieved great success in probabilistic time series imputation tasks with its power to model complex distributions. However, current DDPM-based probabilistic time series imputation methodologies are confronted with two types of challenges: 1)The backbone modules of the denoising parts are not capable of achieving sequence modeling with low time complexity. 2) The architecture of denoising modules can not handle the dependencies in the time series data effectively. To address the first challenge, we explore the potential of state space model, namely Mamba, as the backbone denoising module for DDPMs. To tackle the second challenge, we carefully devise several SSM-based blocks for time series data modeling. Experimental results demonstrate that our approach can achieve state-of-the-art time series imputation results on multiple real-world datasets. Our datasets and code are available at https://github.com/decisionintelligence/SSD-TS/ 11footnotetext: School of Data Science and Engineering. 22footnotetext: Engineering Research Center of Blockchain Data Management, Ministry of Education.
Diffusion Models, State Space Models, Time Series Imputation
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3711896.3737135conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1454-2/2025/08ccs: Mathematics of computing Time series analysisccs: Computing methodologies Neural networks
KDD Availability Link:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.15528129.
1. Introduction
@@
The analysis of time series can model the intrinsic patterns within time-series data, thus providing robust support for decision-making in various fields, such as meteorology (McGovern et al., 2011; Karevan and Suykens, 2020; Miao et al., 2024), financial analysis (Xiang et al., 2022; Owusu et al., 2023; Bai et al., 2020; Li et al., 2025), healthcare (Morid et al., 2023; Poyraz and Marttinen, 2023; Ouyang et al., 2025), power systems (Tzelepi et al., 2023; Zhou et al., 2021; Zhang et al., 2025; Zhao et al., 2023; Wu et al., 2023) and traffic (Guo et al., 2014; Xu et al., 2025, 2023; Liu et al., 2025; Tian et al., 2025). To enhance the reliability of analytical outcomes, it is critical to ensure the integrity of time series. However, due to various reasons such as device failures, human errors, and privacy protection, time series data can easily be incomplete with missing observations at different timestamps.
@@
Time series imputation methods aim to estimate the values of missing points based on the observed points in incomplete time series, thereby restoring the integrity of the time series while preserving its original statistical properties. According to the ability to provide uncertainty of estimations, time series imputation methods can be categorized into the following two perspectives: 1) Deterministic (Cao et al., 2018; Cini et al., 2022; Du et al., 2023), and 2)  Probabilistic (Chen et al., 2023; Kim et al., 2023; Luo et al., 2018) imputation methods. Probabilistic time series imputation is particularly important in dealing with complex and uncertain data environments, as it provides a quantification of uncertainty for the imputations. The key to probabilistic imputation lies in modeling the posterior distribution. Existing probabilistic time series imputation methods include Gaussian Process and Variational Autoencoder-based methods (Fortuin et al., 2020), Normalization Flow-based methods (Rasul et al., 2021b), and Diffusion-based methods (Tashiro et al., 2021). Among these, the Diffusion-based method has emerged as the optimal choice for probabilistic time series due to their accuracy in posterior modeling and adaptability to different scenarios and various types of time series data.
@@
Table 1. Comparison of our method and existing methods in modeling dependencies and time complexity.
Backbone Model CNN Transformer SSM SSD-TS (Ours)
Global Dependency Local Global Partial Global
Time Complexity $\mathcal{O}(L)$ $\mathcal{O}(L^{2})$ $\mathcal{O}(L)$ $\mathcal{O}(L)$
Channel Dependecny Independent Independent Independent Dependent
Inter-sequence Dependency Unidirectional Unidirectional Unidirectional Bidirectional
When selecting a denoising backbone in the diffusion model, the following two key factors need to be considered: 1) Model compatibility, and 2) Time complexity. Model compatibility involves two key aspects: 1) the backbone of the model should be capable of handling input data effectively. 2) the backbone of the model should align with the model’s intended objective (i.e., in diffusion models, the backbone must be capable of modeling noise in the diffusion process). Specifically, the missing observations in time series have correlations with their neighbors on both sides, so it is crucial to design a model by considering information from neighbors of both sides. Moreover, it is also essential to accurately capture the properties of time series, such as global dependencies and channel correlations. Three mainstream denoising backbones are widely used in diffusion models for time series imputation: 1) Convolutional Neural Networks (CNNs)-based, 2) Transformer-based and 3) State-Space Model (SSM)-based backbones. Given a time series with a length of $L$, the CNNs-based backbone can capture partial information from the neighbors within the receptive fields and has $\mathcal{O}(L)$ time complexity. The transformer-based backbone can model temporal dependencies across the entire time series but is with quadratic time complexity $\mathcal{O}(L^{2})$. The SSM backbone has a linear time complexity, $\mathcal{O}(L)$, but it falls short in capturing the information from one side of the neighbor. Moreover, all these backbones fail to capture the channel dependencies in time series. The comparison results of existing backbones and our method in terms of various dependencies and time complexity are presented in Table.1.
Recently, Mamba (Dao and Gu, 2024), a linear sequence modeling approach based on state-space models has demonstrated superior performance in various time series tasks, including time series forecasting, imputation and time series foundation models (Wang et al., 2025; Ma et al., 2024). Compared to Transformer-based models, Mamba-based models have shown stronger sequence modeling capabilities and better performance in time series tasks. However, there has not been research investigating the Mamba model with diffusion models for time series applications. Whether the Mamba model can serve as the backbone of a diffusion model and achieve competitive performance remains an open question.
In this paper, we investigate how to apply Mamba as backbones for time series diffusion models. As shown in Table.1, existing backbones exhibit issues in sequence dependency modeling, time complexity, and inter-channel dependency modeling. To address time complexity, we adopt Mamba of linear complexity, as the foundation for our base module. To tackle intra-sequence and bidirectional dependency modeling, we propose the Bidirectional Attention Mamba (BAM) block, a bidirectional module with temporal attention based on the linear state-space model Mamba, achieving effective and efficient intra-channel dependency modeling. Furthermore, for inter-channel dependencies, we analyze the inductive biases inherent in inter-channel relationships and compare how these biases are captured by the Mamba-based channel modeling module versus CNN- and transformer-based approaches. Our findings indicate that the Mamba-based module provides a more accurate and comprehensive representation, making it a superior choice. Based on this analysis, we design the Channel Mamba Block (CMB) to effectively model inter-channel dependencies.
@@
  1. (1)
    We propose SSD-TS, a State Space Diffusion model for Time Series imputation. It integrates mamba-based blocks as diffusion backbones and equips the model with the capability of probabilistic time series imputation with linear time and space complexity.
  2. (2)
    We give a brief analysis of the characteristics and inductive biases of both intra-channel and inter-channel dependencies, thereby empirically elucidating the superiority of employing mamba modules as a backbone for diffusion models. Based on these analyses, we propose bidirectional attention mamba (BAM) block and channel mamba block (CMB) to achieve effective modeling of these dependencies.
  3. (3)
    We conduct experiments on multiple real-world datasets for time series imputation task. Our approach achieves state-of-the-art performance across several datasets, different missing scenarios and missing ratios, which demonstrate the effectiveness of our proposed method.
The rest of our paper is organized as follows: In Section 2, we present preliminaries of state space models and diffusion models as well as the problem formulation. In Section 3, we give a brief introduction about the diffusion models and the architecture details about our model. In Section 4, we present and analyze the experiment results. In Section 5, we summarize the related work and in Section 6, we conclude our paper.
2. Preliminaries
2.1. State Space Models
State Space Models (SSMs) are an emerging approach to model sequential data, which is implemented by finding out state representations to model the relationship between input and output sequences. An SSM receives a one-dimensional sequence $X\in\mathbb{R}^{L}$ as the input and outputs a corresponding sequence $Y\in\mathbb{R}^{M}$. Under continuous settings, the SSMs are defined according to Eq.1:
(1) $\begin{cases}\dot{h}(t)&=\bm{A}h(t)+\bm{B}x(t)\\ y(t)&=\bm{C}h(t)+\bm{D}x(t),\end{cases}$
@@ 2.1. State Space Models
In real-world applications, the input sequences are discrete samplings of continuous sequences. According to  (Gu et al., 2022), under discrete settings, by applying the zero-order hold technique to Eq.1, it can be reformulated as follows.
(2) $\begin{cases}h_{k}=\bm{\bar{A}}h_{k-1}+\bm{\bar{B}}x_{k}\\ y_{k}=\bm{C}h_{k}\end{cases},$
where $\bm{\bar{A}}=\exp(\Delta\bm{A}),\bm{\bar{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot(\Delta\bm{B})$ and $\Delta$ is the learnable step size in discrete sampling. We can see from Eq.2 that the hidden state is updated according to the input $x(t)$ and last hidden state h(t1)h(t-1) while the output is generated by the hidden state $h(t)$ and the input $x(t)$ and in (Gu et al., 2020), where it introduces High-order Polynomial Projection Operator (Hippo) to achieve longer sequence modeling.
However, it is worth noticing that $\bm{A},\bm{B},\bm{C},\bm{D}$ in Eq.1 and Eq.2 are time-invariant parameters, i.e., they are data-independent parameters and do not change over time. Therefore the model is not capable of assigning different weights at different positions in the input sequence while receiving new inputs. To address this issue, (Gu and Dao, 2023) proposed Mamba, in which the parameter matrices $\bm{A},\bm{B},\bm{C},\bm{D}$ are input-dependent, thus enhancing the performance of sequence modeling. To tackle the problem of non-parallelization, (Gu and Dao, 2023) also introduced selective scan mechanism for effective computing. For further performance and efficiency improvements, (Dao and Gu, 2024) point out that SSMs can be categorized as a variant of linear attention model. In this work, we follow the same architecture of parallel Mamba Blocks as (Dao and Gu, 2024) and a RMS-norm (Zhang and Sennrich, 2019) module is added after the parallel Mamba block. The details of the post-normalization Mamba Block (PNM Block) are illustrated in Fig.3a.
2.2. Diffusion Models
@@
(3) $p_{\theta}(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),$
@@
(4) $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}(x_{t},t)I),$
The loss function of DDPM aims at minimizing the difference between the noise $\epsilon$ in the forward process and the parameterized noise $\epsilon_{\theta}$ in the reverse process:
(5) $\mathcal{L}_{d}=\mathbb{E}_{x_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|,$
where $t$ stands for the diffusion time embedding and $x_{t}$ is calculated in the forward process.
2.3. Problem Formulation
Definition 2.1 (Time Series).
@@ 2.3. Problem Formulation
[…]
@@
where $\bar{X}\in\mathbb{R}^{K\times L}$ is the imputed time series, $\bar{X}_{i,j}=\mu_{i,j}\pm\sigma_{i,j}$ denotes the probabilistic output if $M_{i,j}=1$, otherwise $\bar{X}_{i,j}={X}_{i,j}$.
3. Methodology
Figure 1. The self-supervised framework and training process of SSD-TS. First, some of observed values are masked following the same missing pattern as the missing values (in red) to get masked targets ($X_{0}$, in magenta) and the condition input ($X_{o}^{c}$, in blue). The noisy input is obtain from $X_{0}$ and $\epsilon$ (in orange) sampled from $\mathcal{N}(0,I)$. The objective of the network is to minimize the difference between the parameterized noise $\epsilon_{\theta}(X_{t},t)$ and $\epsilon$. Solid lines in each time series represent observed values, while dashed lines represent missing values.
3.1. Overall Model Architecture
Fig.1 illustrates the overall self-supervised framework and training process of our model. We first mask part of the observed values according to the pattern of missing values, where the masked values serve as the imputation target $X_{0}$ during training. The remaining observed values form the conditional input $X_{o}^{c}$ for the noise prediction network $\epsilon_{\theta}$. We then combine $X_{0}$ with noise $\epsilon$ sampled from a standard normal distribution to obtain the noisy input $X_{t}$. Both $X_{o}^{c}$, $X_{t}$, and the diffusion step $t$ are fed into the noise prediction network $\epsilon_{\theta}$ to get the parameterized noise.
Figure 2. Architecture of noise estimation module $\epsilon_{\theta}$ in SSD-TS.
Fig.2 presents the denoising architecture of our module. The forward process of $\epsilon_{\theta}$ are as follows: For each diffusion step, the input consists of the following parts: noisy input $X_{t}$, the condition input $X_{o}^{c}$ and the diffusion step $t$. To begin with, the inputs are embedded to the latent diffusion space. The embedding module of noisy inputs and condition inputs share a similar model structure, which consists of a linear projection module followed by an SMM block in Fig.3b. The SMM block is composed of stacks of Bidirectional Attention Mamba (BAM) blocks and Channel Mamba Blocks (CMB), which is introduced in the next part. Due to the relatively limited information from $t$, the embedding module of $t$ only consists of linear projection modules. After the embedding step, the embedded diffusion step is concatenated with the input embeddings. The concatenated embeddings are fed in to a SMM module. Then the output of the SMM module is concatenated with the condition embeddings. After feeding the final embeddings to another SMM module and final projection module, we can get the noise predictions $\epsilon_{\theta}(X_{t},t)$.
3.2. Intra-Channel Dependency Modeling
We start by analyzing what the denoising module needs to model within a single channel. Since Triformer (Cirstea et al., 2022) first proposes the Patching technique and existing works (Nie et al., 2023) demonstrate the effectiveness of patching in time series. We do not apply the patching technique to the time series as our approach is non-autoregressive, which means that we generate a complete sequence of size $L$ for a total of $T$ steps ($T<L$). The noise described by the denoising module corresponds to the variations between consecutive steps. This necessitates that the denoising module accurately models the internal dependencies within the sequence in order to precisely predict the noise in the intermediate steps.
As shown in Table.1, the main drawback of convolutional neural network (CNN)-based backbones, such as U-Net (Ronneberger et al., 2015), lies in their inability to capture global information within the sequence. Additionally, stacking convolutional layers increases time complexity. Transformer-based backbones, on the other hand, suffer from a lack of dynamic weight allocation and exhibit quadratic complexity with respect to sequence length $L$. Moreover, existing SSM-based models rely on unidirectional intra-channel dependency modeling, limiting their ability to capture global dependencies in time series data.
To address these issues, we design an intra-channel dependency modeling module based on a bidirectional linear state space model. Specifically, given an input representation $x$, the BAM module first normalizes the input $x$ using layer normalization. The normalized features are processed in both forward and backward directions. The forward processing involves a linear convolutional layer followed by a PNM module. The backward sequence shares a similar processing procedure and the input sequence is flipped along the time dimension. After processing in both directions, a temporal attention module is applied to adjust the weights for each time step. Finally, the results from both directions are element-wise summed after aligning them by position, and a residual connection is added to the original input $x$, producing the output of the BAM module. The detailed structure of the BAM module is shown in Fig.3d.
And it is worth noticing the temporal attention module in Fig.3d is not analogous to the attention mechanism used in Transformer. Our temporal attention module takes a sequence as input and applies a convolutional layer followed by a sigmoid activation to produce a set of weights for each position in the sequence. These weights are then multiplied element-wise with the input sequence. The purpose of this module is to adaptively reweight the temporal positions within the sequence, thereby enhancing the modeling of global dependencies and helping the state space model (SSM) better capture long-range temporal relationships.
3.3. Inter-Channel Dependency Modeling
Addressing the time series modeling problem, capturing inter-channel dependencies allows for the detection of mutual influences among different channels (variables or features), helping to more accurately reveal the global dynamics and complex interactions of the system, thereby improving predictive performance and generalization ability (Qiu et al., 2025b; Wu et al., 2025b).
For more accurate modeling of the channel dependencies, the key is to identify the inductive biases inherent in the dependencies between channels in time series data. Through an analysis of the characteristics of time series data and previous methods (Tashiro et al., 2021; Liu et al., 2023), we argue that the modules designed for modeling channel relationships should possess the capability to describe the following inductive biases:
  • Inter-Channel Dependencies: There are interdependencies among channels and different channels exhibit distinct characteristics.
  • Global Dependencies: There exist long-term interdependencies and co-evolution among the channels.
  • Local Dependencies: Certain groups of channels exhibit stronger internal dependencies compared to channels outside of these groups.
  • Dynamic Dependencies: The dependencies among channels are dynamic. Even for the same time series, the degree of interaction between channels varies on specific downstream tasks.
After identifying the four inductive biases mentioned above, we choose the state space model as the framework for modeling inter-channel dependencies for the following reasons:
  • As shown in Eq.1, the state space model updates through the iterative evolution of the hidden state, representing an adaptive updating mechanism. In contrast to the explicit modeling used in the attention structure, it is more accurate when handling complex, long-term inter-channel interactions. Furthermore, while attention relies on the static attention matrix, its ability to describe dynamic dependencies is inferior to that of the state space models.
  • Convolution channel modeling suffers from the following limitations: On one hand, the description of local dependencies is influenced by the convolution kernel, which assumes that all dependencies are fixed in space (or time) and cannot capture varying local dependencies. On the other hand, the range of dependencies is constrained by the receptive field, making it incapable of directly and effectively modeling global dependencies.
The CMB module takes a latent representation $x\in\mathbb{R}^{B\times C\times L}$ as input. First, the input representation $x$ is transposed to $x^{T}\in\mathbb{R}^{B\times L\times C}$ for processing along the channel dimension, and the transposed representation is passed through a convolution module and a PNM module. In parallel, a combination of a convolutional layer followed by a sigmoid activation is utilized to dynamically adjust the relative relationships between the channels. After a residual connection with the transposed input $x^{T}$, the CMB module transposes the output back to the original shape of $x$ to produce the final output. The details of CMB block are shown in Fig.3c.
Figure 3. Details of PNM, CMB and BAM block in the noise prediction module. (a) PNM: backbone module based on Mamba. (b) SMM: Sequential Mamba module composed of stacks of BAM and CMB. (c) CMB: unidirectional module for inter-channel dependency modeling. (d) BAM: bidirectional module with temporal attention for intra-channel, multi-range dependency modeling.
3.4. Diffusion Models for Time Series Imputation